[The following information applies to the questions displayed below.] Comparative financial statements for Weaver Company follow: Weaver Company Comparative Balance Sheet at December 31 This Year Last Year Assets Cash $ -4 $ 11 Accounts receivable 310 230 Inventory 156 195 Prepaid expenses 9 5 Total current assets 471 441 Property, plant, and equipment 516 437 Less accumulated depreciation (86 ) (72 ) Net property, plant, and equipment 430 365 Long-term investments 25 32 Total assets $ 926 $ 838 Liabilities and Stockholders' Equity Accounts payable $ 304 $ 225 Accrued liabilities 71 79 Income taxes payable 73 65 Total current liabilities 448 369 Bonds payable 196 172 Total liabilities 644 541 Common stock 161 201 Retained earnings 121 96 Total stockholders’ equity 282 297 Total liabilities and stockholders' equity $ 926 $ 838 Weaver Company Income Statement For This Year Ended December 31 Sales $ 753 Cost of goods sold 450 Gross margin 303 Selling and administrative expenses 218 Net operating income 85 Nonoperating items: Gain on sale of investments $ 6 Loss on sale of equipment (2 ) 4 Income before taxes 89 Income taxes 24 Net income $ 65 During this year, Weaver sold some equipment for $18 that had cost $30 and on which there was accumulated depreciation of $10. In addition, the company sold long-term investments for $13 that had cost $7 when purchased several years ago. Weaver paid a cash dividend this year and the company repurchased $40 of its own stock. This year Weaver did not retire any bonds. 2. Using the information in (1) above, along with an analysis of the remaining balance sheet accounts, prepare a statement of cash flows for this year. (List any deduction in cash and cash outflows as negative amounts.)
Answer #1
STATEMENT OF CASH FLOWS
Cash flows from operating activities:
Net income 65
Adjustments to reconcile net income to net cash profived by operations:
Depreciation expense = 86-72+10 = 24
Gain on sale of investments -6
Loss on sale of equipment 2
Increase in accounts receivable = 310-230 = -80
Decrease in inventory = 195-156 = 39
Increase in prepaid expenses =9-5 = -4
Increase in accounts payable = 304-225 = 79
Decrease in accrued liabilities = 79-71 = -8
Increase in income taxes payable = 73-65 = 8 54
Net cash provided by operating activities 119
Cash flows from investing activities:
Purchase of PP&E = 516-437+30 = -109
Sale of long term investments 13
Sale of equipment 18
Net cash used by investing activities -78
Cash flows from financing activities:
Cash from short term note
Sale of bonds =196-172 = 24
Cash Dividends paid = 96+65-121 = -40
Purchase of treasury stock -40
Net cash used by financing activities -56
Net increase (decrease) in cash -15
Beginning cash and cash equivalents 11
Ending cash and cash equivalents -4